I just borrowed $30,000 which I will have to pay off in 5 years together with interest. Interest accumulates at an effective rate of 10.33%. First calculate how much in total I will need to pay off this loan when it comes due, and then based on that answer calculate how much I should deposit into a sinking fund each month to be able to pay off this loan.

Assume that the sinking fund pays 3.68% compounded monthly.
Answer #1

Rate compounded monthly=((1+10.33%)^(1/12)-1)*12=9.87095%

Amount due at the end of the term=30000*(1+9.87095%/12)^(12*5)=49044.40125

Amount to be deposited each month=49044.40125*(3.68%/12)/((1+3.68%/12)^(12*5)-1)=745.7593431